OpenStudy (anonymous):

write the equation in point slope form. is perpemdicular to y= - 3/4 x +4 and passes through point (2, -3)??

OpenStudy (anonymous):

@kelliegirl33 ??

OpenStudy (anonymous):

mind if i helped?

OpenStudy (anonymous):

i dont care??

OpenStudy (anonymous):

you take it Mousam

OpenStudy (anonymous):

oh ok... i thought since he tagged you, you were gonna help him...

OpenStudy (anonymous):

broderick.....If I were you, I wouldn't be rude to people, such as jim_thompson...he has many friends

OpenStudy (anonymous):

hes all yours Mousam

OpenStudy (anonymous):

I guess the first thing to notice is that the equation y=(-3/4)x +4 has the gradient (-3/4). An interesting feature of perpendicular lines is that the multiple of the gradients of both lines equals to -1. Knowing that we can calculate the gradient of this prependicular line by doing the following (-3/4)x=-1 x=(4/3) this means that the gradient of a line perpendicular to y=(-3/4)x +4 has the gradient of (4/3)

OpenStudy (anonymous):

now that we know the gradient of the perpendicular line, we can use the equation (y-y1) = m(x-x1) to get the equation of a line with gradient 'm' that goes throuhg the point (x1,y1)

OpenStudy (anonymous):

since we are trying to find out the equation of the perpendicular line, we can subsitute the gradient and the point that it goes through in to the equation to get y-(-3) = (4/3)(x-2) y+3=(4/3)x - (8/3) y=(4/3)x-(17/3)

OpenStudy (anonymous):

that is it... the equation of the perpendicular line is y=(4/3)x-(17/3)

OpenStudy (anonymous):

is that answer in point slope lol @Mousam

OpenStudy (anonymous):

it is the standard form for writing any linear graph. y=mx+c

OpenStudy (anonymous):

you can multiply both left hand side and right and side by 3 to get 3y=4x-17 (if that helps...)

OpenStudy (anonymous):

point slope is y - y1 = m(x - x1)

OpenStudy (anonymous):

@kelliegirl33 do you have to simplify it from then?

OpenStudy (anonymous):

cause i did..

OpenStudy (anonymous):

yeah @Mousam it has to be in point slope!! we are trying to find the eqaution of the line

OpenStudy (anonymous):

i have done it in point slope...

OpenStudy (anonymous):

if you look at my comment with 3 lines of equation in the end... the point slope form is the second-last line...

OpenStudy (anonymous):

no...you already gave the answer....you just did more then you had to do...point slope for this is : y -(-3) = 4/3(x - 2) or y + 3 = 4/3(x - 2)

OpenStudy (anonymous):

oh... i apologize... in my country, we dont give equation of a linear graph in y-y1=m(x-x1) form. We give it in y=mx+c form...

OpenStudy (anonymous):

you did good Mousam :)

OpenStudy (anonymous):

For that, I sincerely apologize...

OpenStudy (anonymous):

you have no need to apologize

OpenStudy (anonymous):

good job @Mousam i gave you a medal for all the help you gave! job well done! and you to @kelliegirl33! :)

OpenStudy (anonymous):

thank you broderick....just try not to be rude to people.....that will get you nowhere